OpenStudy (anonymous):

A party rental company has chairs and tables for rent. The total cost to rent 12 chairs and 2 tables is $43 . The total cost to rent 3 chairs and 5 tables is $40 . What is the cost to rent each chair and each table?

OpenStudy (shamim):

Let the cost of 1 chair=x

OpenStudy (shamim):

Cost of 1 table=y

OpenStudy (shamim):

So cost of 12 chairs =12x

OpenStudy (shamim):

Cost of 2 table=2y

OpenStudy (shamim):

So we hv first equation 12x+2y=43.........(1) Right?

OpenStudy (anonymous):

and no one answeres

OpenStudy (shamim):

In the same way we can write another equation as 3x+5y=40...........(2)

OpenStudy (anonymous):

you are thinking why did i did this all

OpenStudy (anonymous):

bratgrl88 is not even here

OpenStudy (shamim):

I think he will c these later

OpenStudy (shamim):

Now just solve these 2 equations nd got x=? y=?

OpenStudy (anonymous):

o i see

OpenStudy (shamim):

Frm which country u r ? @damirax

OpenStudy (shamim):

Just for curiosity

OpenStudy (anonymous):

Hi! Thank you! So sorry, it was late my time when I submitted and I decided to hold off on the question in my homework until later. I am in the US. Thank you very much for helping me!! It made so much more sense than the book.